An ideal battery sends a current of 5 A in a resistor. When another resistor of value 10 Ω is connected in parallel, the current through the battery is increased to 6 A. Find the resistance of the first resistor.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. R base 1 = R, i base 1 = 5 A R base 2 = 10R/10 + R, i base 2 = 6A Since potential constant, i base 1R base 1 = i base 2R base 2 ⇒ 5 * R = 6 * 10R/10 + R ⇒ (10 + R)5 = 60 ⇒ 5R = 10 ⇒ R = 2 Ω